# Learning Multimodal Embeddings for Traffic Accident Prediction and Causal Estimation

### v2

@@ Abstract.
We consider analyzing traffic accident patterns using both road network data and satellite images aligned to road graph nodes. Previous work for predicting accident occurrences relies primarily on road network structural features while overlooking physical and environmental information from the road surface and its surroundings. In this work, we construct a large multimodal dataset across six U.S. states, containing nine million traffic accident records from official sources, and one million high-resolution satellite images for each node of the road network. Additionally, every node is annotated with features such as the region’s weather statistics and road type (e.g., residential vs. motorway), and each edge is annotated with traffic volume information (i.e., Average Annual Daily Traffic). Utilizing this dataset, we conduct a comprehensive evaluation of multimodal learning methods that integrate both visual and network embeddings. Our findings show that integrating both data modalities improves prediction accuracy, achieving an average AUROC of $90.1\%$, which is a $3.7\%$ gain over graph neural network models that only utilize graph structures. With the improved embeddings, we conduct a causal analysis based on a matching estimator to estimate the key contributing factors influencing traffic accidents. We find that accident rates rise by $24\%$ under higher precipitation, by $22\%$ on higher-speed roads such as motorways, and by $29\%$ due to seasonal patterns, after adjusting for other confounding factors. Ablation studies confirm that satellite imagery features are essential for achieving accurate prediction. We release the dataset and the experimental code for using this dataset at https://github.com/VirtuosoResearch/MMTraCE.
@@ 1. Introduction
Road safety remains a persistent and pressing challenge in urban cities and districts. According to the World Health Organization, approximately 1.19 million people die in road traffic crashes each year, making it the leading cause of death for children and young adults (undefaab, ). In the United States alone, traffic accidents are projected to cause 39,345 deaths in 2024 and \$871 billion in annual societal costs (undefaaag, ; undefav, ). Faced with such a severe reality, it is imperative to shift from reactive analysis to proactive prediction, using data-driven methods to identify high-risk road areas and implement interventions for accident prevention.
@@
Despite this potential, the use of satellite imagery in state-level traffic accident analysis remains limited. There are two main challenges. The first challenge is the lack of large public multimodal datasets with enough temporal coverage across different regions (undefat, ; undefas, ; undefaaab, ). This gap comes from the heterogeneity of data sources. Each state publishes accident and traffic records in its own format. They also use different variable names and hosting structures. Unifying these datasets requires careful preprocessing and normalization. This includes schema alignment and geospatial matching. The second challenge is the need to reduce the dimensionality of high-resolution satellite images. These images must also be integrated with graph-structural features and other low-dimensional inputs. Examples include weather, traffic volume, and speed limits. Early experiments show that simple fusion methods do not work well. Direct concatenation fails to capture nonlinear interactions between different modalities. As a result, models struggle to use both visual and non-visual information in an effective way.
To address these challenges, we begin by constructing a large-scale, multimodal dataset spanning six U.S. states. It comprises over nine million accident records and one million satellite images with rich feature annotations.
@@
  • Every road is also associated with historical accident occurrences. Notably, we collect the latest accident data up to the year $2025$, with a maximum period covering a temporal span of up to $24$ years, allowing models to capture long-term trends and evolving risk patterns. This dense spatial alignment enables models to jointly learn from both visual cues and graph-structural features.
@@
Building on this dataset, we develop a multimodal learning framework. The framework combines vision-based encoders with GNNs and supports joint reasoning over visual, structural, and contextual properties of the road environment. We conduct a detailed evaluation of different fusion strategies to measure their impact on traffic accident prediction. By adding satellite imagery to the model, we obtain an average AUROC of $90.1\%$ across six states. The prediction accuracy improves by $3.7\%$ compared to GNNs that use only graph-structural features. Based on this result, we further perform a causal analysis of key environmental and traffic factors. We estimate the average treatment effect on the treated group (ATT) using a matching estimator built on the multimodal embeddings. The ATT scores for seasonal variation, road type, and precipitation are $28.6\%$, $21.9\%$, and $24.2\%$. We further estimate the effects with propensity score-based adjustments (PSM) and the double robust estimator (DR). We run a leave-one-out ablation study to measure the contribution of each feature type. Removing image features leads to a drop of $3.5\%$ AUROC. Excluding weather, traffic volume, and road network features results in decreases of $1.8\%$, $2.4\%$, and $3.7\%$, respectively.
In summary, this paper makes three contributions to the study of traffic accidents on road networks. First, we build a large multimodal dataset with recent traffic accident records from six U.S. states. Each satellite image is aligned with a specific road network node. To our knowledge, this is the most comprehensive dataset for evaluating multimodal learning in the transportation domain. Second, we provide a detailed evaluation of multimodal learning methods that combine visual embeddings with network embeddings. The results show clear gains in accident prediction accuracy. Third, we perform a causal analysis to measure the effect of key factors that influence accident occurrences. We use the learned embeddings to control for confounding variables such as road conditions. We hope that this dataset can support future research on multimodal learning for transportation and road safety.
@@ 2. Methodology
We now describe our multimodal dataset and the analytical tools along with the dataset. First, we provide an overview of the problem setting. We then introduce the collection of our dataset, including the Satellite images. Next, we describe the multimodal approaches that integrate both network embeddings and visual embeddings from the images. Finally, we design a causal analysis framework on top of the multimodal embeddings.
@@ 2.1. Preliminaries
Each node is associated with dynamic features that evolve over time, such as weather conditions, including temperature, wind speed, precipitation, and sea-level pressure. Additionally, each node is linked to a high-resolution satellite image centered at its geographic coordinates. We extract visual features that capture the road structure and the surrounding environment from the image.
@@ 2.2. Dataset Collection
The dataset is assembled in several stages. We begin by constructing road network graphs for six states, followed by the collection of high-resolution satellite images that are spatially aligned with the road graph. Finally, we collect historical traffic accident records from official sources and align them with both the road network and satellite images.
@@
Road network. For each state, we construct a detailed road network graph based on data obtained from OpenStreetMap (OSM). [^1] Notice that the original data from OSM divides one edge into several parts. Instead, we have preprocessed the raw data to combine the parts into a single edge. The volume statistic is low due to the large number of residential roads in the data. See Figure 2 for an illustration of the distribution of different types of roads. The road network encompasses five major categories of roadways: city streets, county roads, neighborhood streets, tract roads, and urbanized area roads, ensuring comprehensive spatial coverage across diverse geographic and administrative regions.
Each road segment in the network is represented as a directed edge defined by a start node and an end node, and is annotated with metadata including the road name, whether the segment is one-way, which road type it is, and the physical length of the segment in meters. This representation captures both the structural and functional aspects of the road infrastructure.
@@
Figure 2. The proportion of different road types among six states’ road networks. Residential roads account for the vast majority of the total, making up approximately $74.5\%$ of all roads. Other types, such as tertiary, secondary, and primary, contribute much smaller proportions by comparison.
@@
Satellite image data. Building upon the extracted road network, we obtain high-resolution satellite images centered at each node, each of which typically corresponds to a road intersection. The geographic coordinates of these nodes guide the image collection process, ensuring tight spatial alignment between the visual and structural representations. For each state, we collect hundreds of thousands of satellite images, resulting in a large-scale visual dataset. We implement two strategies for image acquisition: (1) directly fetching static images using the Mapbox Static API, and (2) reconstructing large images by stitching together $25$ smaller tiles. Each final image has a resolution of $1024\times 1024$ pixels and covers approximately $200\,\mathrm{m}\times 200\,\mathrm{m}$ of physical space, capturing rich visual cues and fine-grained road features. In summary, for each network node in the road network, we align one satellite image to that node, for all six states. Therefore, the number of satellite images is equal to the number of nodes.
Accident records. As accident occurrence serves as the primary prediction objective in our framework, we obtain accident records from official sources, specifically the Departments of Transportation (DOT) of each U.S. state, to ensure high reliability and real-world relevance. We collected accident records spanning up to the past two decades, with states providing data ranging from the early 2000s to as recent as 2025. However, these records vary significantly across states in terms of file format, schema, field names, spatial encoding, and level of detail. To enable unified modeling and large-scale analysis, we perform extensive preprocessing and normalization to convert all state-specific datasets into a standardized format. Further, we align their axes with the satellite images.
The overall statistics of our dataset are shown in Table 2, including the period of accidents, total number of accidents, and total number of satellite images. More details about the dataset collection process can be found in Appendix A.
@@ 2.3. Learning Multimodal Embeddings
Next, we use a vision model such as Vision Transformer (ViT) to extract semantic information from the images and encode each image into a visual embedding.
@@
  • Mixture of experts: we leverage multiple specialized networks, referred to as experts, that each learn to combine features from different perspectives. We utilize the gated network that computes the probability distribution over experts for each node.
@@ 2.4. Causal Estimation Using Multimodal Embeddings
where $\hat{y}_{i,1}$ and $\hat{y}_{i,0}$ denote the potential outcomes under treatment and control. Since only one of these outcomes is observed, we approximate the missing counterfactuals through nearest neighbor matching in the embedding space. For each treated sample, we identify its nearest neighbor $j$ from the control set,
@@
This embedding-based approach allows us to control for confounding factors encoded in the multimodal representation. It lets us compare treatment effects under similar road networks and weather conditions. To strengthen this analysis, we also estimate treatment effects with propensity score matching (PSM), which aligns treated and control edges with similar treatment probabilities. In addition, we use a doubly robust (DR) estimator that combines outcome prediction with propensity weighting to provide a more stable estimate. Full details of both procedures are reported in Appendix B.
@@ 3.1.1. Baselines
Second, we utilize common GNNs architectures including GraphSAGE, GCN, GIN, and Graph Transformer (Graphormer) with attention (undefaw, ). We use a spatial-temporal framework such as DCRNN (undefay, ). We also consider supervised contrastive learning to improve the classification of positive and negative edges.
Then, we consider three approaches to integrate graph-based and vision-based features. The first approach basically employs a three-layer multilayer perceptron that jointly processes the concatenated graph and image representations. The second approach uses a gated fusion network that adaptively learns the contribution weights of each modality during training. The last one utilizes the MoE architecture to fetch features from different perspectives.
@@ 3.1.2. Implementations
For each state, we partition the available accident records into training, validation, and test sets based on temporal splits. Specifically, historical data up to a designated cutoff year is used for model training, while accidents occurring after that year are reserved for evaluation. We focus on monthly accident prediction. This avoids future information leakage. We consider both classification and regression tasks. For classification, we report the Area Under the Receiver Operating Characteristic curve (AUROC). For regression, we use the Mean Absolute Error (MAE) to measure the difference between the predicted and observed accident counts on each road segment. We provide additional implementation details, including concrete hyperparameters, formal fusion method definitions, and formal metric definitions in Appendix B.1. The experimental details, including the omitted regression and classification task prediction result, the precision score, and the recall score, are shown in Appendix B.2.
@@ 3.2. Traffic Accident Prediction Results
For GIN-based models, visual features also provide consistent benefits. GIN plus Gated Fusion improves over the vanilla GIN by an average of $2.3\%$ AUROC, with a maximum gain of $6.7\%$. GIN plus MoE improves over the base GIN by an average of $3.6\%$, with the largest gain of $7.9\%$. Taken together, these results show that satellite imagery improves performance for all architectures, and that the relative gains are larger for models with lower base expressivity.
@@
For the feature ablation, we perform a leave-one-out (LOO) analysis on GCN plus Gated Fusion. We consider four feature categories: vision structure, weather observations, road layout, and traffic volumes. In each trial, we remove one category and measure the resulting change in AUROC. The results in Table 3 show that dropping image features leads to a $3.5\%$ decrease. This confirms their strong predictive value. Removing weather features reduces AUROC by $1.8\%$. Excluding traffic volume produces a $2.4\%$ drop. Eliminating road network features causes the largest decrease of $3.7\%$. This highlights the central role of road layout. Adding speed limit as an extra feature yields a small but consistent improvement suggesting that it provides complementary information.
@@
Figure 4. Cross-state AUROC performance of the GIN + MoE model, computed over six states. Each entry shows the score when training on one state (represented by rows) and testing on another state (represented by columns). Darker colors indicate better transferability.
@@
Cross-state transfer. To evaluate how well the model transfers to new regions, we conduct a cross-state generation experiment. In this setting, we train the GIN plus MoE model on one state and test it on another. The AUROC scores for all train–test pairs are shown in Figure 4. The results reveal strong cross-state consistency. We find that the accident patterns in Maryland and Nevada are more consistent with the shared representations learned by the model. It also indicates that both structural and visual features in these two states align well with those from other regions. Further understanding these transfer patterns is an interesting question for future work (undefaaf, ; undefaas, ; undefaaai, ).
Runtime. The fusion structure may introduce extra computation, so we remove one GNN layer in the fusion modules to keep the cost low. We then empirically measure the computational time for evaluating one month of data in each state. The results are reported in Table 4. It shows that the added overhead is small, and the runtime remains almost unchanged across all states.

### v3

@@ Abstract.
We consider analyzing traffic accident patterns using both road network data and satellite images aligned to road graph nodes. Previous work for predicting accident occurrences relies primarily on road network structural features while overlooking physical and environmental information from the road surface and its surroundings. In this work, we construct a large multimodal dataset spanning six U.S. states, containing nine million traffic accident records from official sources, and one million high-resolution satellite images for each node of the road network. Additionally, every node is annotated with features such as the region’s weather statistics and road type (e.g., residential vs. motorway), and each edge is annotated with traffic volume information (i.e., Average Annual Daily Traffic). Utilizing this dataset, we conduct a comprehensive evaluation of multimodal learning methods that integrate both visual and network embeddings. Our findings show that integrating both data modalities improves prediction accuracy, achieving an average AUROC of $90.1\%$, a $3.7\%$ gain over graph neural network models that use only graph structures. With the improved embeddings, we conduct a causal analysis using a matching estimator to identify the key factors influencing traffic accidents. We find that accident rates rise by $24\%$ under higher precipitation, by $22\%$ on higher-speed roads such as motorways, and by $29\%$ due to seasonal patterns, after adjusting for other confounding factors. Ablation studies confirm that satellite imagery features are essential for achieving accurate prediction. We release the dataset and the experimental code for using this dataset at https://github.com/VirtuosoResearch/MMTraCE.
@@ 1. Introduction
Road safety remains a persistent and pressing challenge in urban cities and districts. According to the World Health Organization, approximately 1.19 million people die in road traffic crashes each year, making it the leading cause of death for children and young adults (undefaab, ). In the United States alone, traffic accidents are projected to cause 39,345 deaths in 2024 and \$871 billion in annual societal costs (undefaaag, ; undefav, ). Faced with such a severe reality, it is imperative to shift from reactive analysis to proactive prediction, using data-driven methods to identify high-risk road segments and implement interventions to prevent accidents.
@@
Despite this potential, the use of satellite imagery in state-level traffic accident analysis remains limited. There are two main challenges. The first challenge is the lack of large public multimodal datasets with enough temporal coverage across different regions (undefat, ; undefas, ; undefaaab, ). This gap comes from the heterogeneity of data sources. Each state publishes accident and traffic records in its own format. They also use different variable names and hosting structures. Unifying these datasets requires careful preprocessing and normalization. This includes schema alignment and geospatial matching. The second challenge is the need to reduce the dimensionality of high-resolution satellite images. These images must also be integrated with graph-structural features and other low-dimensional inputs. Examples include weather, traffic volume, and speed limits. Early experiments show that simple fusion methods do not work well. Direct concatenation fails to capture nonlinear interactions between different modalities. As a result, models struggle to effectively use both visual and non-visual information.
To address these challenges, we begin by constructing a large-scale, multimodal dataset spanning six U.S. states. It comprises over $9$ million accident records and $1$ million satellite images, each with rich feature annotations.
@@
  • Every road is also associated with historical accident occurrences. Notably, we collect the latest accident data through $2025$, with a maximum temporal span of $24$ years, allowing models to capture long-term trends and evolving risk patterns. This dense spatial alignment enables models to jointly learn from both visual cues and graph-structural features.
@@
Building on this dataset, we develop a multimodal learning framework. The framework combines vision-based encoders with GNNs and supports joint reasoning over visual, structural, and contextual properties of the road environment. We conduct a detailed evaluation of different fusion strategies to measure their impact on traffic accident prediction. By adding satellite imagery to the model, we obtain an average AUROC of $90.1\%$ across six states. The prediction accuracy improves by $3.7\%$ compared to GNNs that use only graph-structural features. Based on this result, we further perform a causal analysis of key environmental and traffic factors. We estimate the average treatment effect on the treated (ATT) using a matching estimator based on multimodal embeddings. The ATT scores for seasonal variation, road type, and precipitation are $28.6\%$, $21.9\%$, and $24.2\%$. We further estimate the effects with propensity score-based adjustments (PSM) and the double robust estimator (DR). We run a leave-one-out ablation study to measure the contribution of each feature type. Removing image features leads to a $3.5\%$ drop in AUROC. Excluding weather, traffic volume, and road network features results in decreases of $1.8\%$, $2.4\%$, and $3.7\%$, respectively.
In summary, this paper makes three contributions to the study of traffic accidents on road networks. First, we build a large multimodal dataset with recent traffic accident records from six U.S. states. Each satellite image is aligned with a specific road network node. To our knowledge, this is the most comprehensive dataset for evaluating multimodal learning in the transportation domain. Second, we provide a detailed evaluation of multimodal learning methods that combine visual embeddings with network embeddings. The results show clear gains in accident prediction accuracy. Third, we conduct a causal analysis to measure the effects of key factors influencing accident occurrences. We use the learned embeddings to control for confounding variables such as road conditions. We hope that this dataset can support future research on multimodal learning for transportation and road safety.
@@ 2. Methodology
We now describe our multimodal dataset and the accompanying analytical tools. First, we provide an overview of the problem setting. We then introduce the dataset collection, including the Satellite images. Next, we describe the multimodal approaches that integrate network and visual image embeddings. Finally, we design a causal analysis framework on top of the multimodal embeddings.
@@ 2.1. Preliminaries
Each node is associated with dynamic features that evolve over time, such as weather conditions, including temperature, wind speed, precipitation, and sea-level pressure. Additionally, each node is linked to a high-resolution satellite image centered at its geographic coordinates. We extract visual features from the image that capture the road structure and the surrounding environment.
@@ 2.2. Dataset Collection
The dataset is assembled in several stages. We begin by constructing road network graphs for six states, then collect high-resolution satellite images spatially aligned with these road graphs. Finally, we collect historical traffic accident records from official sources and align them with both the road network and satellite images.
@@
Road network. For each state, we construct a detailed road network graph based on data obtained from OpenStreetMap (OSM). [^1] Notice that the original data from OSM divides one edge into several parts. Instead, we have preprocessed the raw data to combine the parts into a single edge. The volume statistic is low because the data includes many residential roads. See Figure 2 for an illustration of the distribution of different types of roads. The road network encompasses five major categories of roadways: city streets, county roads, neighborhood streets, tract roads, and urbanized area roads, ensuring comprehensive spatial coverage across diverse geographic and administrative regions.
Each road segment in the network is represented as a directed edge defined by a start node and an end node, and is annotated with metadata including the road name, whether the segment is one-way, the road type, and the segment’s physical length in meters. This representation captures both the structural and functional aspects of the road infrastructure.
@@
Figure 2. The proportion of different road types among six states’ road networks. Residential roads account for the vast majority of all roads, making up approximately $74.5\%$. Other types, such as tertiary, secondary, and primary, contribute much smaller proportions by comparison.
@@
Satellite image data. Building upon the extracted road network, we obtain high-resolution satellite images centered at each node, each of which typically corresponds to a road intersection. The geographic coordinates of these nodes guide the image collection process, ensuring tight spatial alignment between the visual and structural representations. For each state, we collect hundreds of thousands of satellite images, resulting in a large-scale visual dataset. We implement two strategies for image acquisition: (1) directly fetching static images using the Mapbox Static API, and (2) reconstructing large images by stitching together $25$ smaller tiles. Each final image has a resolution of $1024\times 1024$ pixels and covers approximately $200\,\mathrm{m}\times 200\,\mathrm{m}$ of physical space, capturing rich visual cues and fine-grained road features. In summary, for each node in the road network, we align one satellite image to that node across all six states. Therefore, the number of satellite images equals the number of nodes.
Accident records. As accident occurrence serves as the primary prediction objective in our framework, we obtain accident records from official sources, specifically the Departments of Transportation (DOT) of each U.S. state, to ensure high reliability and real-world relevance. We collected accident records dating back up to two decades, with states providing data ranging from the early 2000s to 2025. However, these records vary significantly across states in terms of file format, schema, field names, spatial encoding, and level of detail. To enable unified modeling and large-scale analysis, we perform extensive preprocessing and normalization to convert all state-specific datasets into a standardized format. Further, we align their axes with the satellite images.
The overall statistics of our dataset are shown in Table 2, including the accident period, the total number of accidents, and the total number of satellite images. More details about the dataset collection process can be found in Appendix A.
@@ 2.3. Learning Multimodal Embeddings
Next, we use a vision model, such as Vision Transformer (ViT), to extract semantic information from images and encode each image into a visual embedding.
@@
  • Mixture of experts: we leverage multiple specialized networks, referred to as experts, each of which learns to combine features from different perspectives. We utilize the gated network that computes the probability distribution over experts for each node.
@@ 2.4. Causal Estimation Using Multimodal Embeddings
where $\hat{y}_{i,1}$ and $\hat{y}_{i,0}$ denote the potential outcomes under treatment and control. Since only one of these outcomes is observed, we approximate the missing counterfactuals using nearest-neighbor matching in the embedding space. For each treated sample, we identify its nearest neighbor $j$ from the control set,
@@
This embedding-based approach allows us to control for confounding factors encoded in the multimodal representation. It lets us compare treatment effects under similar road networks and weather conditions. To strengthen this analysis, we also estimate treatment effects using propensity score matching (PSM), which matches treated and control units with similar treatment probabilities. In addition, we use a doubly robust (DR) estimator that combines outcome prediction with propensity weighting to provide a more stable estimate. Full details of both procedures are reported in Appendix B.
@@ 3.1.1. Baselines
Second, we utilize common GNN architectures, including GraphSAGE, GCN, GIN, and Graph Transformer (Graphormer) with attention (undefaw, ). We use a spatial-temporal framework such as DCRNN (undefay, ). We also consider supervised contrastive learning to improve the classification of positive and negative edges.
We then consider three approaches to integrating graph-based and vision-based features. The first approach basically employs a three-layer multilayer perceptron that jointly processes the concatenated graph and image representations. The second approach uses a gated fusion network that adaptively learns the contribution weights of each modality during training. The last one uses the MoE architecture to extract features from multiple perspectives.
@@ 3.1.2. Implementations
For each state, we partition the available accident records into training, validation, and test sets based on temporal splits. Specifically, historical data up to a designated cutoff year is used to train the model, while accidents occurring after that year are reserved for evaluation. We focus on monthly accident prediction. This avoids future information leakage. We consider both classification and regression tasks. For classification, we report the Area Under the Receiver Operating Characteristic curve (AUROC). For regression, we use the Mean Absolute Error (MAE) to measure the difference between the predicted and observed accident counts on each road segment. We provide additional implementation details, including concrete hyperparameters, formal definitions of the fusion method, and formal definitions of the metrics, in Appendix B.1. The experimental details, including the omitted regression and classification task prediction result, the precision score, and the recall score, are shown in Appendix B.2.
@@ 3.2. Traffic Accident Prediction Results
For GIN-based models, visual features also provide consistent benefits. GIN plus Gated Fusion improves over the vanilla GIN by an average of $2.3\%$ AUROC, with a maximum gain of $6.7\%$. GIN plus MoE improves over the base GIN by an average of $3.6\%$, with the largest gain of $7.9\%$. Taken together, these results show that satellite imagery improves performance across all architectures and that the relative gains are greater for models with lower base expressivity.
@@
For the feature ablation, we perform a leave-one-out (LOO) analysis on GCN plus Gated Fusion. We consider four feature categories: vision structure, weather observations, road layout, and traffic volumes. In each trial, we remove one category and measure the resulting change in AUROC. The results in Table 3 show that dropping image features leads to a $3.5\%$ decrease. This confirms their strong predictive value. Removing weather features reduces AUROC by $1.8\%$. Excluding traffic volume produces a $2.4\%$ drop. Eliminating road network features causes the largest decrease of $3.7\%$. This highlights the central role of road layout. Adding speed limit as an extra feature yields a small but consistent improvement, suggesting that it provides complementary information.
@@
Figure 4. Cross-state AUROC performance of the GIN + MoE model, computed over six states. Each entry shows the score when training on one state (rows) and testing on another (columns). Darker colors indicate better transferability.
@@
Cross-state transfer. To evaluate how well the model transfers to new regions, we conduct a cross-state generalization experiment. In this setting, we train the GIN plus MoE model on one state and test it on another. The AUROC scores for all train–test pairs are shown in Figure 4. The results reveal strong cross-state consistency. We find that the accident patterns in Maryland and Nevada are more consistent with the model’s learned shared representations. It also indicates that both structural and visual features in these two states align well with those from other regions. Further understanding these transfer patterns is an interesting question for future work (undefaaf, ; undefaas, ; undefaaai, ).
Runtime. The fusion structure may introduce extra computation, so we remove one GNN layer in the fusion modules to keep the cost low. We then empirically measure the computational time required to evaluate one month of data in each state. The results are reported in Table 4. It shows that the added overhead is small, and the runtime remains almost unchanged across all states.